OpenStudy (anonymous):

In the system shown below, what is the sum of the x-coordinates of all solutions? 7x^2+3y^2=187 3y^2-7x=47

OpenStudy (mertsj):

3y^2=47+7x from the second equation. Put that in the first equation in place of 3y^2

OpenStudy (mertsj):

7x^2+7x+47=187

OpenStudy (mertsj):

7x^2+7x-140=0 Divide by 7 x^2+x-20=0

OpenStudy (anonymous):

x=20?

OpenStudy (mertsj):

(x+5)(x-4)=0 x=-5 or 4 So the sum of the x coordinates of the solutions is -5+4 or -1

OpenStudy (anonymous):

nevermind i see

OpenStudy (anonymous):

thanks!! :)

OpenStudy (mertsj):

yw